# Event-Grounding Graph: Unified Spatio-Temporal Scene Graph from Robotic Observations

### v1

@@ I Introduction
Roboticists have long sought to create intelligent robots that act as information agents capable of perceiving and understanding their surroundings while performing physical tasks to assist us in our daily lives. Central to their functionality is the representation they build of their surroundings. As mobile robots increasingly become an integral part of our lives, they inevitably collect vast amounts of sensor data over time that is essential to comprehending both the semantic and dynamic nature of their surroundings.
There has been a surge in research aimed at embedding semantic features into spatial scene elements. Most notable are the development of 3D scene graphs [1, 2, 3]. By representing spatial concepts such as objects, freespace, and rooms as nodes and their inter-relationships as edges, 3DSGs boast rich semantic features while remaining compact in memory footprint. While this body of work enriches the representation of the environment, facilitating basic navigation tasks such as object search, it does not adequately capture dynamic events, i.e., interactions that result in changes to the scene, frequently occurring within those environments.
Studies exist that focus on recording and verbalizing robot experiences and observations by leveraging the capabilities of visual-language models [4, 5, 6]. This line of research enhances robots’ understanding of how the scene can be interacted with and enables them to respond to free-form spatio-temporal questions from humans, such as "When did you last see my mug?" or navigation tasks such as "Find me a coffee machine." However, these works describe the environment using free-form natural language captions, which often leads to a lack or misrepresentation of the spatial features of the environment. This insufficiency can lead to issues, such as a lack of multi-view consistency or the aggregation of redundant information when the same environment is revisited multiple times.
@@
In this work, we propose a novel framework, EGG, to aggregate, organize, and describe robotic observations over time, creating a unified representation that includes both spatial features and dynamic events. The rich information from the framework efficiently enables the extraction of relevant information subsets based on specific queries or tasks, ensuring insightful and contextually appropriate responses from robots, as illustrated in Fig. LABEL:fig:cover. E.g., for the same query in the Fig. LABEL:fig:cover, the spatial components could only identify "my mug", and the event components could only retrieve mug-related activities without identifying the mug referred to by the query. We then demonstrate the capability of our representation through a series of experiments using real robotic data. The core contributions of this work are:
@@ II Related works
Another significant body of research focuses on developing memory representations that incorporate interactions with the environment. Bärmann et al. [5] propose a framework that systematically records a robot’s interaction history within a hierarchical tree-like structure. This representation includes raw robotic sensory inputs at its base layer and abstracted descriptions of the robot’s experiences at higher layers. Closer to our work, ReMEmbR [4] constructs a spatio-temporal representation that aggregates video captions generated by a VLM, along with timestamps and the robot’s positional data, into a vector database. Another relevant work is Embodied-RAG [6], which develops a spatio-temporal representation in the form of a semantic forest. This representation is constructed by embedding video captions captured at various points in time into the nodes of a topological graph, with a  large language model (LLM) summarizer generating higher-level nodes from these embeddings. While these approaches acknowledge the importance of interactions with the environment, they exhibit limitations concerning the grounding of recorded events in the spatial features of the environment. This leads to issues such as multi-view inconsistency and the unnecessary accumulation of redundant data, exemplified by the repetitive description of the same scene upon revisit. Our work addresses this gap.
With the rapid advancements of LLMs, semantically enriched representations have emerged as information providers, enabling robots to conduct complex tasks like memory verbalization [5], information retrieval, and semantic navigation [6, 4]. However, as LLMs are easily distracted by redundant information [11], semantic search algorithms are needed to provide these models with task-relevant information from the scene representation. While advancing beyond the state-of-the-art in semantic search is outside of the scope of our work, we demonstrate that EGG is equipped with the needed spatio-temporal queryability to facilitate the development of such methods. For instance, SayPlan [7] serializes 3DSGs into JSON format, incorporating them into LLM inputs to facilitate semantic task planning. The framework utilizes the hierarchical structure of 3D scene graphs to identify a relevant subgraph for each specific task by expanding and contracting the graph through iterative prompting. As a result, the method reduces token usage while ensuring high accuracy in task execution. Similarly, H-EMV [5] employs an LLM agent to iteratively search for a pertinent node in the memory representation, optionally engaging a VLM to extract additional insights from raw data in the base layer. Embodied-RAG [6] utilizes an LLM agent to identify the most relevant top-layer nodes based on semantic context, subsequently proceeding down the hierarchy to retrieve relevant waypoints for executing semantic navigation tasks.
@@ III Problem Formulation
We define an event that occurs in the time interval $T$ as a sequence of interactions executed by either the observer or an external agent within the scene, each resulting in an attribute change of at least one tracked spatial element at a particular instant $t\in T$. We denote an event by $E_{T}=(O_{T},D_{T},s_{T})$. $O_{T}\subseteq S$ is the set of spatial elements that experience attribute changes during the event. Attribute changes, $\mathcal{V}_{T}=\{\mathcal{V}_{T}^{i}\ |\ S^{i}\in O_{T}\}$, may include changes to objects’ positions, e.g., someone moving the chairs, the appearance or disappearance of objects, e.g., someone taking a mug outside of the robot’s view, or modifications in their operational states due to interactions, e.g., someone turning the lights on or off. $D_{T}$ is the set of element-wise descriptions of each change, providing insight into the role that each spatial element plays in the event. Formally, for every attribute change $\mathcal{V}_{T}^{i}\in\mathcal{V}_{T}$, a corresponding description $d^{i}_{T}\in D_{T}$ exists that verbally describes the change. $s_{T}$ is the event summary, which provides an overall description of the event.
@@ IV Methodology
(a) Spatial and event components connections
(b) Grounding an event node $n^{E}_{0:1}$ with a spatial node $n^{S}_{1}$. The event $E_{0:1}$ occurs during the interval $t_{0},t_{1}]$ and changes the position of $S^{1}$.
Figure 1: EGG architecture
@@ IV-A EGG generation
EGG consists of two interrelated sets of components: spatial components and event components, as illustrated in Fig. 1(a). Formally, EGG is the graph $\mathcal{G}=(\mathcal{N}^{S},\mathcal{N}^{E},\mathcal{E}^{S},\mathcal{E}^{E})$, with $\mathcal{N}^{S}$ and $\mathcal{N}^{E}$ being the spatial and event nodes, and $\mathcal{E}^{S}$ and $\mathcal{E}^{E}$ being the spatial and event edges.
The spatial components $\mathcal{G}^{S}=(\mathcal{N}^{S},\mathcal{\mathcal{E}}^{S})$ follow a hierarchical 3DSG structure analogous to that of Hydra [2, 3]. Each spatial node $n^{S}_{i}\in\mathcal{N}^{S}$ corresponds to the attribute set of a spatial element, so $n^{S}_{i}=S^{i}$. However, unlike static scene graph frameworks such as [2, 3, 1], spatial elements in EGG also include time variant attributes, which track changes in spatial elements resulting from event interactions. Each spatial edge $e^{S}_{il}=(T_{il},n^{S}_{i},n^{S}_{l})$ connecting two nodes, $n^{S}_{i}$ and $n^{S}_{l}$, represents spatial relationships such as inclusion, e.g., the mug is in the kitchen. Since nodes are not permanent in EGG, edges are additionally characterized by the time interval $T_{il}$ during which the spatial relationship exists.
The event components $\mathcal{G}^{E}=(\mathcal{N}^{E},\mathcal{E}^{E})$ represent the observed events and their relationships with the interacted spatial elements. Fig. 1(b) illustrates how an event is grounded with a single spatial element. To each event $E_{T}$ corresponds an event node $n^{E}_{T}=(p_{T},s_{T})$ defined spatially by the observation points of the robot $p_{T}$, and temporally by the respective time interval $T$. The event overall description is provided by the summary $s_{T}$. Each event edge $e^{E}_{Ti}=(T,n^{E}_{T},n^{S}_{i},d^{i}_{T})$ connects the event $E_{T}$ to the spatial element $S^{i}\in O_{T}$. Similar to the spatial edges, the event edges are also timestamped and exist only in the interval $T$ when the event occurs. The event description $d^{i}_{T}\in D_{T}$ describes the role of the spatial element $S^{i}$ in the event.
@@ IV-B EGG pruning
To this end, we assume the information relevant to a query, denoted as $I_{Q}$, can be split into four types: time interval $T_{Q}$, locations $L_{Q}$, spatial elements $S_{Q}$, and relevant events $E_{Q}$. We will describe how to retrieve $I_{Q}$ from $Q$ in Sec. V-B. Given the set of relevant information $I_{Q}=(T_{Q},L_{Q},S_{Q},E_{Q})$ and an EGG $\mathcal{G}$, we retrieve the query-relevant subgraph $\mathcal{G}_{Q}\subseteq\mathcal{G}$ by using a series of graph manipulation functions, illustrated in Fig. 2 and detailed below. Each function takes as input a subgraph $\mathcal{G}^{\prime}=\{\mathcal{N}^{S^{\prime}},\mathcal{E}^{S^{\prime}},\mathcal{N}^{E^{\prime}},\mathcal{E}^{E^{\prime}}\}\subseteq\mathcal{G}$ and a subset of $I_{Q}$ to return another subgraph $\mathcal{G}^{*}=\{\mathcal{N}^{S*},\mathcal{E}^{S*},\mathcal{N}^{E*},\mathcal{E}^{E*}\}\subseteq\mathcal{G}$. This is achieved first by selecting the list of relevant nodes $\{\mathcal{N}^{S*},\mathcal{N}^{E*}\}$ according to a set of rules, and then by adding all edges connected to the selected nodes to $\mathcal{G}^{*}$, i.e., $\mathcal{E}^{S*}=\{e^{S}_{ij}\ |\ n^{S}_{i},n^{S}_{j}\in\mathcal{N}^{S*}\}$ and $\mathcal{E}^{E*}=\{e^{E}_{Ti},\ |\ n^{S}_{i}\in\mathcal{N}^{S*},n^{E}_{T}\in\mathcal{N}^{E*}\}$. Specifically, the graph manipulation functions are:
@@
  • Location pruning $F^{L}(\mathcal{G}^{\prime},L_{Q})$: From a list of locations $L_{Q}$, we select the subgraph $\mathcal{G}^{*}$ that consists of events and spatial elements from those locations. Specifically, since the spatial components alone form a hierarchical 3DSG structure, the locations list $L_{Q}\in\mathcal{N}^{S^{\prime}}$ is essentially a list of spatial nodes that exists on a higher layers in the scene graphs, such as buildings, rooms or asset objects (e.g., the table is an asset object node and the chairs nodes are its children). Thus $\mathcal{N}^{S*}=L_{Q}\cup\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{S}_{l}\in L_{Q})\ e^{S}_{li}\in\mathcal{E}^{S^{\prime}}\}$, and $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E^{\prime}}\ |\ (\exists n^{S}_{l}\in L_{Q})\ e^{E}_{Tl}\in\mathcal{E}^{E^{\prime}}\}$.
  • Spatial element pruning $F^{S}(\mathcal{G}^{\prime},S_{Q})$: From a list of query-relevant spatial nodes $S_{Q}\subseteq\mathcal{N}^{S^{\prime}}$, we select the subgraph $\mathcal{G}^{*}$ that contains relevant spatial elements and the events that involve them. Specifically, we prune the graph such that the event nodes are connected to the relevant spatial elements, $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E^{\prime}}\ |\ (\exists n^{S}_{i}\in S_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$. Then, the preserved spatial nodes are the ones connected to at least one of those events, i.e., $\mathcal{N}^{S*}=\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{E}_{T}\in\mathcal{N}^{E*})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$.
  • Event pruning $F^{E}(\mathcal{G}^{\prime},E_{Q})$: From a list of query-relevant event nodes $E_{Q}\subseteq\mathcal{N}^{E^{\prime}}$, we select the subgraph $\mathcal{G}^{*}$ that contains relevant events and the spatial elements that they involve. The event nodes in this case are just $\mathcal{N}^{E*}=E_{Q}$, and the spatial nodes are those that are connected to at least one of these events, i.e., $\mathcal{N}^{S*}=\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{E}_{T}\in E_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$.
  • History expansion $F^{H}(\mathcal{G},\mathcal{G}^{\prime},S_{Q})$: From a list of spatial elements $S_{Q}\subseteq\mathcal{N}^{S^{\prime}}$, we expand the graph to include the history of events that involve them. The event nodes that involve $S_{Q}$ are $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E}\ |\ (\exists n^{S}_{i}\in S_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E}\}$. The spatial nodes include those that are given by $S_{Q}$, and related to the expanded event history, i.e., $\mathcal{N}^{S*}=S_{Q}\cup\{n^{S}_{i}\in\mathcal{N}^{S}\ |\ (\exists n^{E}_{T}\in\mathcal{N}^{E*})\ e^{E}_{Ti}\in\mathcal{E}^{E}\}$.
@@
$\mathcal{G}_{Q}=\mathcal{C}(F^{T},F^{L},F^{S},F^{E},F^{H})(\mathcal{G},I_{Q})$ (1)

### v2

@@ I Introduction
Memory representations that encapsulate both the geometric and semantic aspects of the environment are crucial for robots to perceive and understand their surroundings while performing physical tasks to assist us in our daily lives. For instance, to answer a question "Where is my phone?", a robot needs to complete three steps: identify objects that could be classified as "phones," determine which one is specifically "my phone," and then locate it. Humans typically recall an object’s position based on previous experiences, such as where they last used it or where they commonly use it.
There has been a surge in research aimed at embedding semantic features into spatial scene elements. Most notable are the development of 3D scene graphs [2, 7, 11, 12, 27, 9]. By representing spatial concepts such as objects, freespace, and rooms as nodes and their inter-relationships as edges, 3DSGs boast rich semantic features while remaining compact in memory footprint. While this body of work incorporates semantic information into geometric features of the representation, e.g., grounding the concept of "my phone" into a node, they do not capture dynamic events, i.e., interactions that result in changes to the scene, frequently occurring within those environments.
Another line of research focuses on recording and verbalizing robot experiences and observations by leveraging the capabilities of visual-language models [1, 4, 29]. This line of research enhances robots’ understanding of how the scene can be interacted with and enables them to respond to free-form spatio-temporal questions from humans. However, these works often describe the environment using free-form natural language captions, which often leads to a lack or misrepresentation of the spatial features of the environment, e.g., multiple events of "using a phone" can be captured as text captions, but they do not provide information about which "phone" the person is using.
@@
In this work, we propose a novel framework, EGG, to aggregate, organize, and describe robotic observations over time, creating a unified representation that includes both spatial features and dynamic events. The rich information from the framework efficiently enables the extraction of relevant information subsets based on specific queries or tasks, ensuring insightful and contextually appropriate responses from robots, as illustrated in Fig. LABEL:fig:cover. We then demonstrate the capability of our representation through a series of experiments using real robotic data. The core contributions of this work are:
@@ II Related works
Another significant body of research focuses on developing memory representations that incorporate interactions with the environment. Bärmann et al. [3] propose a framework that systematically records a robot’s interaction history within a hierarchical tree-like structure. This representation includes raw robotic sensory inputs at its base layer and abstracted descriptions of the robot’s experiences at higher layers. Closer to our work, ReMEmbR [1] constructs a spatio-temporal representation that aggregates video captions generated by a VLM, along with timestamps and the robot’s positional data, into a vector database. Another relevant work is Embodied-RAG [29], which develops a spatio-temporal representation in the form of a semantic forest. This representation is constructed by embedding video captions captured at various points in time into the nodes of a topological graph, with a LLM summarizer generating higher-level nodes from these embeddings. While these approaches acknowledge the importance of interactions with the environment, they exhibit limitations concerning the grounding of recorded events in the spatial features of the environment. This leads to issues such as multi-view inconsistency and the unnecessary accumulation of redundant data, exemplified by the repetitive description of the same scene upon revisit. Our work addresses this gap.
@@ III Problem Formulation
We define an event that occurs in the time interval $T$ as a sequence of interactions executed by either the observer or an external agent within the scene, each resulting in an attribute change of at least one tracked spatial element at a particular instant $t\in T$. We denote an event by $E_{T}=(O_{T},D_{T},s_{T})$. $O_{T}\subseteq S$ is the set of spatial elements that experience attribute changes during the event. Attribute changes, $\mathcal{V}_{T}=\{\mathcal{V}_{T}^{i}\ |\ S^{i}\in O_{T}\}$, may include changes to objects’ positions, e.g., someone moving the chairs, the appearance or disappearance of objects, e.g., someone taking a mug outside of the robot’s view, or modifications in their operational states due to interactions, e.g., someone turning the lights on or off. $D_{T}$ is the set of element-wise descriptions of each change, providing insight into the role that each spatial element plays in the event. Formally, for every attribute change $\mathcal{V}_{T}^{i}\in\mathcal{V}_{T}$, a corresponding description $d^{i}_{T}\in D_{T}$ exists that verbally describes the change. $s_{T}$ is the event summary, which provides an overall description of the event. Collectively, an event $E_{T}$ not only outlines the scene change, but also describes how each spatial element is affected during the interval $T$.
@@ IV Methodology
Figure 1: Spatial and event components connections. Spatial and event edges are not permanent and are characterized by time intervals.
@@ IV-A EGG generation
EGG consists of two interrelated sets of components: spatial components and event components, as illustrated in Fig. 1. Formally, EGG is the graph $\mathcal{G}=(\mathcal{N}^{S},\mathcal{N}^{E},\mathcal{E}^{S},\mathcal{E}^{E})$, with $\mathcal{N}^{S}$ and $\mathcal{N}^{E}$ being the spatial and event nodes, and $\mathcal{E}^{S}$ and $\mathcal{E}^{E}$ being the spatial and event edges.
The spatial components $\mathcal{G}^{S}=(\mathcal{N}^{S},\mathcal{\mathcal{E}}^{S})$ follow a hierarchical 3DSG structure analogous to that of Hydra [11, 12]. Each spatial node $n^{S}_{i}\in\mathcal{N}^{S}$ corresponds to the attribute set of a spatial element, so $n^{S}_{i}=S^{i}$. In this work, the constructed 3DSG consists of two layers: rooms and objects. However, the formulation from Sec. III does not prevent the inclusion of other spatial concepts, such as buildings and free space [11, 12] or asset objects [5]. Unlike static scene graph frameworks such as [11, 12, 7], objects in EGG also include time variant attributes, which track changes in objects resulting from event interactions. Each spatial edge $e^{S}_{il}=(T_{il},n^{S}_{i},n^{S}_{l})$ connecting two nodes, $n^{S}_{i}$ and $n^{S}_{l}$, represents spatial relationships such as inclusion, e.g., the mug is in the kitchen. Since nodes are not permanent in EGG, edges are additionally characterized by the time interval $T_{il}$ during which the spatial relationship exists.
The event components $\mathcal{G}^{E}=(\mathcal{N}^{E},\mathcal{E}^{E})$ represent the observed events and their relationships with the interacted objects. Fig. 1 illustrates how events are grounded with multiple objects. To each event $E_{T}$ corresponds an event node $n^{E}_{T}=(p_{T},s_{T})$ defined spatially by the observation points of the robot $p_{T}$, i.e., the sensor’s pose, and temporally by the respective time interval $T$. The event overall description is provided by the summary $s_{T}$. Each event edge $e^{E}_{Ti}=(T,n^{E}_{T},n^{S}_{i},d^{i}_{T})$ connects the event $E_{T}$ to the object $S^{i}\in O_{T}$. Similar to the spatial edges, the event edges are also timestamped and exist only in the interval $T$ when the event occurs. The event description $d^{i}_{T}\in D_{T}$ describes the role of the object $S^{i}$ in the event.
@@ IV-B EGG pruning
To this end, we assume the information relevant to a query, denoted as $I_{Q}$, can be split into four types: time interval $T_{Q}$, locations $L_{Q}$, objects $O_{Q}$, and events $E_{Q}$. We will describe how to retrieve $I_{Q}$ from $Q$ in Sec. V-B. Given the set of relevant information $I_{Q}=(T_{Q},L_{Q},O_{Q},E_{Q})$ and an EGG $\mathcal{G}$, we retrieve the query-relevant subgraph $\mathcal{G}_{Q}\subseteq\mathcal{G}$ by using a series of graph manipulation functions, illustrated in Fig. 2 and detailed below. Each function takes as input a subgraph $\mathcal{G}^{\prime}=\{\mathcal{N}^{S^{\prime}},\mathcal{E}^{S^{\prime}},\mathcal{N}^{E^{\prime}},\mathcal{E}^{E^{\prime}}\}\subseteq\mathcal{G}$ and a subset of $I_{Q}$ to return another subgraph $\mathcal{G}^{*}=\{\mathcal{N}^{S*},\mathcal{E}^{S*},\mathcal{N}^{E*},\mathcal{E}^{E*}\}\subseteq\mathcal{G}$. This is achieved first by selecting the list of relevant nodes $\{\mathcal{N}^{S*},\mathcal{N}^{E*}\}$ according to a set of rules, and then by adding all edges connected to the selected nodes to $\mathcal{G}^{*}$, i.e., $\mathcal{E}^{S*}=\{e^{S}_{ij}\ |\ n^{S}_{i},n^{S}_{j}\in\mathcal{N}^{S*}\}$ and $\mathcal{E}^{E*}=\{e^{E}_{Ti},\ |\ n^{S}_{i}\in\mathcal{N}^{S*},n^{E}_{T}\in\mathcal{N}^{E*}\}$. Specifically, the graph manipulation functions are:
@@
  • Location pruning $F^{L}(\mathcal{G}^{\prime},L_{Q})$: From a list of locations $L_{Q}$, we select the subgraph $\mathcal{G}^{*}$ that consists of events and objects from those locations. Specifically for this work, the locations list $L_{Q}\in\mathcal{N}^{S^{\prime}}$ is essentially a list of spatial nodes that exists on the room layer. Thus $\mathcal{N}^{S*}=L_{Q}\cup\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{S}_{l}\in L_{Q})\ e^{S}_{li}\in\mathcal{E}^{S^{\prime}}\}$, and $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E^{\prime}}\ |\ (\exists n^{S}_{l}\in L_{Q})\ e^{E}_{Tl}\in\mathcal{E}^{E^{\prime}}\}$.
  • Object pruning $F^{O}(\mathcal{G}^{\prime},O_{Q})$: From a list of query-relevant object nodes $O_{Q}\subseteq\mathcal{N}^{S^{\prime}}$, we select the subgraph $\mathcal{G}^{*}$ that contains relevant objects and the events that involve them. Specifically, we prune the graph such that the event nodes are connected to the relevant objects, $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E^{\prime}}\ |\ (\exists n^{S}_{i}\in O_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$. Then, the preserved spatial nodes are the ones connected to at least one of those events, i.e., $\mathcal{N}^{S*}=\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{E}_{T}\in\mathcal{N}^{E*})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$.
  • Event pruning $F^{E}(\mathcal{G}^{\prime},E_{Q})$: From a list of query-relevant event nodes $E_{Q}\subseteq\mathcal{N}^{E^{\prime}}$, we select the subgraph $\mathcal{G}^{*}$ that contains relevant events and the objects that they involve. The event nodes in this case are just $\mathcal{N}^{E*}=E_{Q}$, and the spatial nodes are those that are connected to at least one of these events, i.e., $\mathcal{N}^{S*}=\{n^{S}_{i}\in\mathcal{N}^{S^{\prime}}\ |\ (\exists n^{E}_{T}\in E_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E^{\prime}}\}$.
  • History expansion $F^{H}(\mathcal{G},\mathcal{G}^{\prime},O_{Q})$: From a list of objects $O_{Q}\subseteq\mathcal{N}^{S^{\prime}}$, we expand the graph to include the history of events that involve them. The event nodes that involve $O_{Q}$ are $\mathcal{N}^{E*}=\{n^{E}_{T}\in\mathcal{N}^{E}\ |\ (\exists n^{S}_{i}\in O_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E}\}$. The spatial nodes include those that are given by $O_{Q}$, and related to the expanded event history, i.e., $\mathcal{N}^{S*}=O_{Q}\cup\{n^{S}_{i}\in\mathcal{N}^{S}\ |\ (\exists n^{E}_{T}\in\mathcal{N}^{E*})\ e^{E}_{Ti}\in\mathcal{E}^{E}\}$.
@@
$\mathcal{G}_{Q}=\mathcal{C}(F^{T},F^{L},F^{O},F^{E},F^{H})(\mathcal{G},I_{Q})$ (1)